# Learning Distributions on Manifolds with Free-form Flows

### v1

@@ Abstract
Many real world data, particularly in the natural sciences and computer vision, lie on known Riemannian manifolds such as spheres, tori or the group of rotation matrices. The predominant approaches to learning a distribution on such a manifold require solving a differential equation in order to sample from the model and evaluate densities. The resulting sampling times are slowed down by a high number of function evaluations. In this work, we propose an alternative approach which only requires a single function evaluation followed by a projection to the manifold. Training is achieved by an adaptation of the recently proposed free-form flow framework to Riemannian manifolds. The central idea is to estimate the gradient of the negative log-likelihood via a trace evaluated in the tangent space. We evaluate our method on various manifolds, and find significantly faster inference at competitive performance compared to previous work. We make our code public at https://github.com/vislearn/FFF.Preprint. Under review.
@@ 1 Introduction
Figure 1: Manifold free-form flows (M-FFF) learn a distribution directly on the manifold via a generator network in the embedding space $\tilde{g}_{\phi}:\mathbb{R}^{m}\to\mathbb{R}^{m}$ (top left) whose outputs are projected to the manifold (top right). In this example, a uniform distribution (bottom left) is mapped to a bimodal distribution on the circle (bottom right), which we show as a histogram over model samples and the learned density as a function of the angle. The embedding conserves the topology so that the resulting distribution is periodic and continuous everywhere.
In recent years, generative models have achieved remarkable success in various domains like image synthesis [44], natural language processing [6], science [39] and more. However, the approaches are not naively applicable when dealing with data inherently structured in non-Euclidean spaces, particularly prevalent in fields such as the natural sciences and computer vision. Examples of such data are earth science data supported on a sphere, angles between functional groups in proteins, which are naturally represented on a torus, as well as the orientation of real world objects given as a rotation matrix in $SO(3)$. Naively, data on such an $n$-dimensional manifold $\mathcal{M}$ can be represented using internal coordinates such as angles. However, this may not respect the topology, causing discontinuities or artifacts.
Luckily, many generative models can be generalized to arbitrary manifolds. Most existing methods involve solving differential equations, either stochastic (SDEs) or ordinary (ODEs), for sampling and density estimation such as [45, 30, 20, 10, 8]. While theoretically sound, these approaches are computationally intensive due to the need for numerous function evaluations during integration, slowing down sampling.
In this paper, we introduce a novel approach for modeling distributions on arbitrary Riemannian manifolds that circumvents the computational burden of previous methods. We achieve this by using a single feed-forward neural network on an embedding space as a generator whose outputs are projected to the manifold Fig. 1. We learn this network as a normalizing flow, made possible through adapting the free-form flow framework [11], recently proposed for Euclidean spaces, to Riemannian manifolds. The core change is estimating the gradient of the negative log-likelihood within the tangent space of the manifold.
@@
  • We adapt free-form flows to Riemannian manifolds, yielding manifold free-form flows (M-FFF) in Sec. 4. The model is trained end-to-end and involves only a single function evaluation in training and sampling.
  • We demonstrate the model’s quality and speed on several manifold benchmarks, such as rotation matrices, earth data, and molecular torsion angles in Sec. 5. We find competitive quality to recent work at typically two orders of magnitude faster sampling.
Together, manifold free-form flows are a novel approach for learning distributions on manifolds. They are applicable to any Riemannian manifold with a known embedding and projection.
@@ 2 Related work
Existing work on learning distributions on manifolds can be broadly categorized as follows: (i) leveraging Euclidean generative models; (ii) building specialized architectures that respect one particular kind of geometry; and (iii) learning a continuous time process on the manifold. We compare our method to these approaches in Tab. 1 and give additional detail below.
Euclidean generative models. One approach maps the $n$-dimensional manifold to $\mathbb{R}^{n}$ and learns the resulting distribution [17]. Another approach generalizes the reparameterization trick to Lie groups by sampling on the Lie algebra which can be parameterized in Euclidean space [16]. These approaches come with the downside that a Euclidean representation may not respect the geometry of the manifold sufficiently, e.g. mapping the earth to a plane causes discontinuities at the boundaries. This can be overcome by learning distributions on overlapping charts that together span the full manifold [23]. An orthogonal solution is to embed the data and add noise to it in the off-manifold directions, so that the distribution can be learnt directly in an embedding space $\mathbb{R}^{m}$ [5]; this only gives access to an ELBO instead of the exact density. Our method also works in the embedding space so that it respects the geometry of the manifold, but directly optimizes the likelihood on the manifold.
Specialized architectures take advantage of the specific geometry of a certain kind of manifold to come up with special coupling blocks for building normalizing flows such as $SO(3)$ [27], $SU(n),U(n)$ [3, 24]; hyperbolic space [2]; and tori and spheres [43]. Manifold free-form flows are not restricted to one particular manifold, but can be easily applied to any manifold for which an embedding and a projection to the manifold is known. As such, our model is an alternative to all of the above specialized architectures.
Continuous time models build a generative model based on parameterizing a ODE or SDE on any Riemannian manifold, meaning that they specify the (stochastic) differential equation in the tangent space [45, 14, 15, 20, 30, 10, 8, 28, 1]. These methods come with the disadvantage that sampling and density evaluation integrates the ODE or SDE, requiring many function evaluations. Our manifold free-form flows do not require repeatedly evaluating the model, a single function call followed by a projection is sufficient.
At its core, our method adapts the recently introduced free-form flow (FFF) framework [11] based on an estimator for the gradient of the change of variables formula [46]. We give more details in Sec. 3.1.
Respect topology Single step sampling Arbitrary manifolds
@@
Table 1: Comparison of generative models on manifolds. We give a “” if any method in a category meets this requirement.
@@ 3 Background
In this section, we give some background for our method: We present an introduction to free-form flows, followed by Riemannian manifolds and distributions thereon.
@@ 3.1 Free-form flows
Free-form flows are a generative model that modify normalizing flows to work with arbitrary feed-forward neural network architectures [11]. Traditionally, normalizing flows are based on invertible neural networks (INNs) that learn an invertible transformation $z=f_{\theta}(x)$ from data $x\in\mathbb{R}^{n}$ to latent codes $z\in\mathbb{R}^{n}$. This gives an explicit parameterized probability density $p_{\theta}(x)$ via the change-of-variables:
@@
where $f_{\theta}^{\prime}(x)\in\mathbb{R}^{n\times n}$ is the Jacobian matrix of $f_{\theta}(x)$ with respect to $x$, evaluated at $x$; $|f_{\theta}^{\prime}(x)|$ is its absolute determinant. The distribution of latent codes $p_{Z}(z)$ is chosen such that the log-density is easy to evaluate and it is easy to sample from. Normalizing flows can be trained by minimizing the negative log-likeliriemannianhood over the training data:
$\min_{\theta}\mathcal{L}_{\text{NLL}}=\min_{\theta}\mathbb{E}_{x\sim p_{\text{data}}}[\log p_{\theta}(x)].$ (2)
@@
This is equivalent to minimizing the Kullback-Leibler-divergence between the true data distribution and the parameterized distribution $\operatorname{KL}(p_{\text{data}}\|p_{\theta})$. Sampling from the model is achieved by pushing samples from the latent distribution $z\sim p_{Z}$ through the inverse of the learned function: $x=f_{\theta}^{-1}(z)\sim p_{\theta}(x)$.
@@
Free-form flows (FFF, [11]) soften the restriction that the learned model be invertible. Instead, they learn two separate networks, an encoder $f_{\theta}$ and a decoder $g_{\phi}$ coupled by a reconstruction loss:
$\mathcal{L}_{\text{R}}=\mathbb{E}_{q(x)}\left[\|g_{\phi}(f_{\theta}(x))-x\|^{2}\right].$ (3)
Encoder and decoder replace $f_{\theta}$ respectively its inverse $f_{\theta}^{-1}$ of an invertible neural network.
Typical normalizing flow architectures are optimized to allow efficient computation of the change of variables term $\log|f_{\theta}^{\prime}(x)|$. This usually goes hand in hand with a constrained architecture that is analytically and practically invertible. Free-form flows are based on arbitrary neural networks and the assumption of a tractable change of variables does not hold in general. FFFs therefore estimate the gradient of the change of variables via the gradient trick proposed by [46]:
$\nabla_{\theta}\mathcal{L}_{\text{NLL}}$ $=\nabla_{\theta}\mathbb{E}_{q(x)}\left[-\log p_{Z}(z)-\operatorname{tr}(f_{\theta}^{\prime}(x)\texttt{SG}[g_{\phi}^{\prime}(z)])\right]$ (4)
$=\nabla_{\theta}\mathbb{E}_{x,v}\left[-\log p_{Z}(z)-v^{T}f_{\theta}^{\prime}(x)\texttt{SG}[g_{\phi}^{\prime}(z)v]\right]$ (5)
Here, $z=f_{\theta}(x)$, $v\in\mathbb{R}^{n}$ is a random vector with unit covariance, and SG (stop-grad) ensures that gradients only flow to the encoder parameters $\theta$. Instead of $\mathcal{O}(n)$ automatic differentiation steps to compute the full Jacobian $f_{\theta}^{\prime}(x)$, this estimator only requires a single vector-Jacobian and a Jacobian-vector product, each of which require one automatic differentiation operation. Note that while the gradient in Eq. 5 is meaningful, the value of $\mathcal{L}_{\text{NLL}}$ is not.
@@
In Sec. 4, we demonstrate how this loss can be used to estimate the gradient of the change of variables formula on manifolds, which we now come to.
@@ 3.2 Riemannian manifolds
A manifold is a fundamental concept in mathematics, providing a framework for describing and analyzing spaces that locally resemble Euclidean space, but may have different global structure. For example, a small region on a sphere is similar to Euclidean space, but walking in a straight line on the sphere in any direction will return back to the starting point, unlike in Euclidean space.
Mathematically, an $n$-dimensional manifold, denoted as $\mathcal{M}$, is a space where every point has a neighborhood that is topologically equivalent to $\mathbb{R}^{n}$. A Riemannian manifold $(\mathcal{M},G)$ extends the concept of a manifold by adding a Riemannian metric $G$ which introduces a notion of distances and angles. At each point $x$ on the manifold, there is an associated tangent space $\mathcal{T}_{x}\mathcal{M}$ which is an $n$-dimensional Euclidean space, characterizing the directions in which you can travel and still stay on the manifold. The metric $G$ acts in this space, defining an inner product between vectors. From this inner product, we can compute the length of paths along the manifold and hence distances between two points.
In this paper, we consider Riemannian manifolds globally embedded into an $m$-dimensional Euclidean space $\mathbb{R}^{m}$, with $n\leq m$. Embedding means that we represent a point on the manifold $x\in\mathcal{M}$ as a vector in $\mathbb{R}^{m}$ confined to an $n$-dimensional subspace. A global embedding is a smooth, injective mapping of the entire manifold into a $\mathbb{R}^{m}$, its smoothness preserving the topology.
We only work with isometrically embedded manifolds, meaning that the metric is inherited from the ambient Euclidean space. Intuitively, this means that the length of a path on the manifold is just the Euclidean length of that path. We note that due to the Nash embedding theorem [36], every Riemannian manifold has a smooth isometric embedding into Euclidean space of some finite dimension, so in this sense using isometric embeddings is not a limitation. Nevertheless, for some manifolds (especially with negative curvature, e.g. hyperbolic space) there may not be a sensible isometric embedding. We provide a generalization of our method to such manifolds in Sec. 7.2 but do not pursue it experimentally. In Tab. 2, we list the manifolds considered in this paper, their dimension, embedding and how to project from the embedding space to the manifold.
@@
3.3 Change of variables on manifolds
Figure 2: Changing variables between manifolds requires calculations in the tangent spaces of $\mathcal{M}_{X}$ and $\mathcal{M}_{Z}$. Specifically, when changing variables from $x$ to $f(x)$, we need to account for the volume change of the transformation between $\mathcal{T}_{x}\mathcal{M}_{X}$ and $\mathcal{T}_{f(x)}\mathcal{M}_{Z}$. Consider a basis $\{e_{i}\}$ for the tangent space at $x$. This basis is transformed under $f$ into $\{e_{i}^{\prime}\}$ where $e_{i}^{\prime}=f^{\prime}(x)e_{i}$. The volume change is the ratio between the areas of the parallelograms (purple regions) formed by $e_{i}$ and $e_{i}^{\prime}$, equal to the determinant of $f^{\prime}(x)$, with an additional factor if the manifold is not isometrically embedded, see Sec. 7.1 for details.
Let $(\mathcal{M}_{X},G_{X})$ and $(\mathcal{M}_{Z},G_{Z})$ be $n$-dimensional Riemannian manifolds. Let $f:\mathcal{M}_{X}\rightarrow\mathcal{M}_{Z}$ Then the probability density induced by changing variables with $f$ is
$\log p_{X}(x)=\log p_{Z}(f(x))+\log|f^{\prime}(x)|+\tfrac{1}{2}\log\frac{|G_{Z}(f(x))|}{|G_{X}(x)|}$ (7)
with $f^{\prime}(x)$ the Jacobian of $f$ at $x$ and ||$\cdot$|| ⋅ | the absolute determinant. Some care has to be taken with the determinant. Strictly speaking, $f^{\prime}$ is a linear map between $\mathcal{T}_{x}\mathcal{M}_{X}$ and $\mathcal{T}_{f(x)}\mathcal{M}_{Z}$ and the determinant should be evaluated in those spaces, as visualized in Fig. 2. Naively evaluating in $\mathbb{R}^{m}$ will give zero if $n<m$: see Sec. 7.1 for a more rigorous treatment and a derivation of Eq. 7.
For the remainder of this paper, we will drop the index on the manifolds $\mathcal{M}_{X}$ and $\mathcal{M}_{Z}$ and simply write $\mathcal{M}$. We do this to simplify notation, but also we use $\mathcal{M}_{X}=\mathcal{M}_{Z}$ in all cases. Note that our method easily generalizes to using different manifolds in data and latent space, as long as they are topologically equivalent.
Along with the simplification that the manifolds are isometrically embedded, meaning $|G_{X}|=|G_{Z}|=1$, we have
$\log p_{X}(x)=\log p_{Z}(f(x))+\log|f^{\prime}(x)|,$ (8)
which is identical to the familiar change of variables formula used in normalizing flows, with the determinant being evaluated in the tangent spaces.

### v2

@@ Abstract
We propose Manifold Free-Form Flows (M-FFF), a simple new generative model for data on manifolds. The existing approaches to learning a distribution on arbitrary manifolds are expensive at inference time, since sampling requires solving a differential equation. Our method overcomes this limitation by sampling in a single function evaluation. The key innovation is to optimize a neural network via maximum likelihood on the manifold, possible by adapting the free-form flow framework to Riemannian manifolds. M-FFF is straightforwardly adapted to any manifold with a known projection. It consistently matches or outperforms previous single-step methods specialized to specific manifolds, and is competitive with multi-step methods with typically two orders of magnitude faster inference speed. We make our code public at https://github.com/vislearn/FFF.
@@ 1 Introduction
Generative models have achieved remarkable success in various domains such as image synthesis (Rombach et al., 2022), natural language processing (Brown et al., 2020), scientific applications (Noé et al., 2019) and more. However, the approaches are not directly applicable when dealing with data inherently structured in non-Euclidean spaces, which is common in fields such as the natural sciences, computer vision, and robotics. Examples include earth science data on a sphere, the orientation of real-world objects given as a rotation matrix in $SO(3)$, or data on special geometries modeled by meshes or signed distance functions. Representing such data naively using internal coordinates, such as angles, can lead to topological issues, causing discontinuities or artifacts.
Luckily, many generative models can be adapted to handle data on arbitrary manifolds. However, the predominant methods compatible with arbitrary Riemannian manifolds involve solving differential equations—stochastic (SDEs) or ordinary (ODEs)—for sampling and density estimation (Rozen et al., 2021; Mathieu and Nickel, 2020; Huang et al., 2022; De Bortoli et al., 2022; Chen and Lipman, 2024). These methods are computationally intensive due to the need for numerous function evaluations during integration, slowing down inference.
Figure 1: Manifold Free-Form Flows (M-FFF) learn generative models on a variety of manifolds. (Left) The learned distributions (colored surface) accurately match the test points (black dots). (Right) We parameterize M-FFF using a neural network in an embedding space, whose outputs are projected to the manifold. This enables simulation-free training and inference, and naturally respects the corresponding geometry, yielding fast sampling and continuous distributions regardless of the manifold.
To address these challenges, we introduce a novel approach for modeling distributions on arbitrary Riemannian manifolds that circumvents the computational burden of previous methods. This is achieved by using a single feed-forward neural network on an embedding space as a generator, with outputs projected to the manifold (Fig. 1). We learn this network as a normalizing flow, facilitated by generalizing the free-form flow framework (Draxler et al., 2024; Sorrenson et al., 2024) to Riemannian manifolds. The core innovation is estimating the gradient of the negative log-likelihood within the tangent space of the manifold.
@@
  • We extend free-form flows to Riemannian manifolds, yielding manifold free-form flows (M-FFF) in Section 4.
  • M-FFF can easily be adapted to arbitrary Riemannian manifolds, requiring only a projection function from an embedding space.
  • It only relies on a single function evaluation during training and sampling, speeding up inference over multi-step methods typically by two orders of magnitude.
  • M-FFF consistently matches or outperforms previous single-step methods on several benchmarks on spheres, tori, rotation matrices, hyperbolic space and curved surfaces (see Fig. 1 and Section 5). In addition, it is competitive with multi-step methods.
Together, manifold free-form flows offer a novel and efficient approach for learning distributions on manifolds, applicable to any Riemannian manifold with a known embedding and projection.
@@ 2 Related work
Table 1: Feature comparison of generative models on manifolds. We give a “” if any method in a category meets this requirement.
Respects topology Single step sampling Arbitrary manifolds
@@
Existing work on learning distributions on manifolds can be broadly categorized as follows: (i) leveraging Euclidean generative models; (ii) building specialized architectures that respect one particular kind of geometry; and (iii) learning a continuous time process on the manifold. We compare our method to these approaches in Table 1 and give additional detail below.
Euclidean generative models. One approach maps the $n$-dimensional manifold to $\mathbb{R}^{n}$ and learns the resulting distribution (Gemici et al., 2016). Another approach generalizes the reparameterization trick to Lie groups by sampling on the Lie algebra which can be parameterized in Euclidean space (Falorsi et al., 2019). These approaches come with the downside that a Euclidean representation may not respect the geometry of the manifold sufficiently, e.g. mapping the earth to a plane causes discontinuities at the boundaries. This can be overcome by learning distributions on overlapping charts that together span the full manifold (Kalatzis et al., 2021). An orthogonal solution is to embed the data and add noise to it in the off-manifold directions, so that the distribution can be learnt directly in an embedding space $\mathbb{R}^{m}$ (Brofos et al., 2021); this only gives access to an ELBO instead of the exact density. Our method also works in the embedding space so that it respects the geometry of the manifold, but directly optimizes the likelihood on the manifold.
Specialized architectures take advantage of the specific geometry of a certain kind of manifold to come up with special coupling blocks for building normalizing flows such as $SO(3)$ (Liu et al., 2023), $SU(d),U(d)$ (Boyda et al., 2021; Kanwar et al., 2020); hyperbolic space (Bose et al., 2020); tori and spheres (Rezende et al., 2020). Manifold free-form flows are not restricted to one particular manifold, but can be easily applied to any manifold for which an embedding and a projection to the manifold is known. As such, our model is an alternative to all of the above specialized architectures.
Continuous time models build a generative model based on parameterizing an ODE or SDE on any Riemannian manifold, meaning that they specify the (stochastic) differential equation in the tangent space (Rozen et al., 2021; Falorsi, 2021; Falorsi and Forré, 2020; Huang et al., 2022; Mathieu and Nickel, 2020; De Bortoli et al., 2022; Chen and Lipman, 2024; Lou et al., 2020; Ben-Hamu et al., 2022). These methods come with the disadvantage that sampling and density evaluation integrates the ODE or SDE, requiring many function evaluations. Our manifold free-form flows do not require repeatedly evaluating the model, a single function call followed by a projection is sufficient.
At its core, our method generalizes the recently introduced free-form flow (FFF) framework (Draxler et al., 2024) based on an estimator for the gradient of the change of variables formula (Sorrenson et al., 2024). We give more details in Section 3.1.
@@ 3 Background
In this section, we provide the background for our method: We present an introduction to free-form flows and Riemannian manifolds.
@@ 3.1 Free-form flows
Free-form flows are a class of generative models that generalize normalizing flows to work with arbitrary feed-forward neural network architectures (Draxler et al., 2024).
Euclidean change-of-variables
Traditionally, normalizing flows are based on invertible neural networks (INNs, see Kobyzev et al. (2021) for an overview) that learn an invertible transformation $z=f_{\theta}(x)$ mapping from data $x\in\mathbb{R}^{n}$ to latent codes $z\in\mathbb{R}^{n}$. This gives an explicitly parameterized probability density $p_{\theta}(x)$ via the change-of-variables:
@@
where $f_{\theta}^{\prime}(x)\in\mathbb{R}^{n\times n}$ is the Jacobian matrix of $f_{\theta}(x)$ with respect to $x$, evaluated at $x$; $|f_{\theta}^{\prime}(x)|$ is its absolute determinant. The distribution of latent codes $p_{Z}(z)$ is chosen such that the log-density is easy to evaluate and it is easy to sample from, such as a standard normal. Normalizing flows can be trained by minimizing the negative log-likelihood over the training data distribution:
$\min_{\theta}\mathcal{L}_{\text{NLL}}=\min_{\theta}\mathbb{E}_{p_{\text{data}}(x)}[-\log p_{\theta}(x)].$ (2)
@@
This is equivalent to minimizing the Kullback-Leibler-divergence between the true data distribution and the parameterized distribution $\operatorname{KL}(p_{\text{data}}\|p_{\theta})$. Sampling from the model is achieved by pushing samples from the latent distribution $z\sim p_{Z}$ through the inverse of the learned function: $x=f_{\theta}^{-1}(z)\sim p_{\theta}$.
Euclidean gradient estimator
Naively computing the volume change $\log|f_{\theta}^{\prime}(x)|$ in Eq. 1 is expensive since it contains the full Jacobian matrix $f^{\prime}_{\theta}(x)\in\mathbb{R}^{n\times n}$, requiring $\mathcal{O}(n)$ automatic differentiation steps to compute. Normalizing flow architectures usually avoid this expensive computation by further restricting the architecture to allow fast computation. Luckily, even if such a shortcut is not available, its gradient, which is all we need for training, can still be efficiently estimated as follows:
Theorem 1 (Volume change gradient estimator, Draxler et al. (2024)).
Let $f_{\theta}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be a diffeomorphism. Let $v\in\mathbb{R}^{n}$ be a random variable with zero mean and unit covariance. Then, the derivative of the volume change has the following trace expression, where $z=f_{\theta}(x)$:
$\nabla_{\theta}\log|f_{\theta}^{\prime}(x)|$ $=\operatorname{tr}((\nabla_{\theta}f_{\theta}^{\prime}(x))f_{\theta}^{\prime-1}(z))$ (3)
$=\mathbb{E}_{v}\left[v^{T}(\nabla_{\theta}f_{\theta}^{\prime}(x)){f_{\theta}^{-1}}^{\prime}(z)v\right].$ (4)
Replacing the expected value over $v$ by a single sample, and using a stop-grad (SG) operation, Theorem 1 allows us to compute a term whose gradient is an unbiased estimator for the gradient of Eq. 2:
$\nabla_{\theta}\mathcal{L}_{\text{NLL}}(x)\approx\nabla_{\theta}(-\log p_{Z}(z)-v^{T}f_{\theta}^{\prime}(x)\texttt{SG}({f_{\theta}^{-1}}^{\prime}(z)v)).$ (5)
Comparing Eqs. 1 and 5 reveals that $\log|f_{\theta}^{\prime}(x)|$ is replaced by a single vector-Jacobian $v^{T}f_{\theta}^{\prime}(x)$ and a Jacobian-vector product ${f_{\theta}^{-1}}^{\prime}(z)v$, each of which require only one automatic differentiation operation. Note that while the gradient estimate is unbiased, computing the term in the brackets is not informative about $\mathcal{L}_{\text{NLL}}$. Thus, for density estimation at inference, Eq. 1 is explicitly evaluated using the full Jacobian.
Free-form architectures
@@
The central idea of free-form flows is to soften the restriction that the learned model be invertible. Instead, they learn two separate networks, an encoder $f_{\theta}$ and a decoder $g_{\phi}$ coupled by a reconstruction loss, circumventing the need for an invertible neural network $f_{\theta}$:
$\mathcal{L}_{\text{R}}=\mathbb{E}_{p_{\text{data}}(x)}\left[\|g_{\phi}(f_{\theta}(x))-x\|^{2}\right].$ (6)
@@
This allows replacing constrained invertible architectures with free-form neural networks. Since $f_{\theta}$ is not restricted to efficiently compute the volume change, free-form flows use Eq. 5 to compute the gradient of $\mathcal{L}_{\text{NLL}}$. To compute Eq. 5, free-form flows approximate ${f_{\theta}^{-1}}^{\prime}(z)$ (which is not tractable) by $g_{\phi}^{\prime}(z)$ during training:
$\nabla_{\theta}\mathcal{L}_{\text{NLL}}(x)\approx\nabla_{\theta}(-\log p_{Z}(z)-v^{T}f_{\theta}^{\prime}(x)\texttt{SG}(g_{\phi}^{\prime}(z)v)).$ (8)
For density estimation at inference, Draxler et al. (2024) recommend using the explicit decoder Jacobian for the volume change.
@@ 3.2 Riemannian manifolds
A manifold is a fundamental concept in mathematics, providing a framework for describing and analyzing spaces that locally resemble Euclidean space, but may have different global structure. For example, a small region on a sphere is similar to the Euclidean plane, but walking in a straight line on the sphere in any direction will return back to the starting point, unlike on a plane.
Mathematically, an $n$-dimensional manifold, denoted as $\mathcal{M}$, is a space where every point has a neighborhood that is topologically equivalent to $\mathbb{R}^{n}$. A Riemannian manifold $(\mathcal{M},G)$ extends the concept of a manifold by adding a Riemannian metric $G$ which introduces a notion of distances and angles. At each point $x$ on the manifold, there is an associated tangent space $\mathcal{T}_{x}\mathcal{M}$ which is an $n$-dimensional Euclidean space, characterizing the directions in which you can travel and still stay on the manifold. The metric $G$ acts in this space, defining an inner product between vectors. From this inner product, we can compute the length of paths along the manifold, distances between points as well as volumes (see next section).
In this paper, we consider Riemannian manifolds globally embedded into an $m$-dimensional Euclidean space $\mathbb{R}^{m}$, with $n\leq m$. Embedding means that we represent a point on the manifold $x\in\mathcal{M}$ as a vector in $\mathbb{R}^{m}$ confined to an $n$-dimensional subspace; we write $x\in\mathcal{M}\subseteq\mathbb{R}^{m}$ and denote by $\pi:\mathbb{R}^{m}\to\mathcal{M}$ a projection from the embedding space to the manifold. A global embedding is a smooth, injective mapping of the entire manifold into $\mathbb{R}^{m}$, its smoothness preserving the topology.
In most cases, we work with, but are not limited to, isometrically embedded manifolds, meaning that the metric is inherited from the ambient Euclidean space. Intuitively, this means that the length of a path on the manifold is just the length of the path in the embedding space. We note that due to the Nash embedding theorem (Nash, 1956), every Riemannian manifold has a smooth isometric embedding into Euclidean space of some finite dimension, so in this sense using isometric embeddings is not a limitation. Nevertheless, for some manifolds (especially with negative curvature, e.g. hyperbolic space) there may not be a sensible isometric embedding.
@@
Generic $\operatorname{rank}(\pi^{\prime}(\pi(x)))$$\{x\in\mathbb{R}^{m}:\pi(x)=x\}$ $x\mapsto\pi(x)$
@@
Hyperbolic $\mathbb{H}^{n}$ $n$$\{x\in\mathbb{R}^{n}:\|x\|<1\}$ $x\mapsto x\min\{1,(1-\epsilon)/\|x\|\}$